OpenStudy (anonymous):

prove that this is a one to one function. t(x)=2X+cosX

OpenStudy (beginnersmind):

Here's an outline: 1. Show that the limit at \[-\infty\] is \[-\infty\] and at \[\infty\] is \[\infty\] 2. Show that the function is continuous 3, Show that it's strictly increasing along its domain 1 and 2 proves that for any y there exists an x such that f(x)=y 2 and 3 can be used to show that if f(u)=f(v) then u=v.